OpenStudy (anonymous):

need to find an equation of the line having the given slope and containg the given point. need to write in slope-intercept form m=7/8, (3,-1)

OpenStudy (anonymous):

y=mx+c -1=7/8 * 3 + c c= -29/8 so eq is y=7/8*x -29/8

OpenStudy (anonymous):

this is correct..The one above

OpenStudy (anonymous):

Thanks a lot for confirming it =)

OpenStudy (anonymous):

i told u it was right

OpenStudy (anonymous):

yeah you did =)